OpenStudy (anonymous):

simplest form of -6 12/10

OpenStudy (skullpatrol):

$$\Huge -6 \frac{12}{10}$$ $$\Huge=-(6 \frac{12}{10})$$$$\Huge=-(6+ \frac{12}{10}) $$$$\Huge=-(6+ \frac{\cancel{12}6}{\cancel{10}5}) $$ $$\Huge=-(\frac{6}{1}+ \frac{6}{5}) $$ $$\Huge=-(\frac{6\cdot5}{1\cdot5}+ \frac{6}{5}) $$ $$\Huge=-(\frac{30}{5}+ \frac{6}{5}) $$ $$\Huge=-(\frac{30+6}{5}) $$ $$\Huge=-\frac{36}{5} $$